AAPL has expecting earnings of $8 per share and an ROE of 15%. If AAPL has a plowback ratio of 70% and a cost of capital of 12% .

What is AAPL’s current price? and What is AAPL’s present value of growth options (PVGO)?
Answer #1

We see that APPL's current price is given as equal to=Expected EPS*(1-plowback ratio)/(cost of capital-ROE*plowback ratio)=8*(1-70%)/(12%-15%*70%)=160.00

PVGO=Current Price-Expected EPS/cost of capital=160.00-8/12%=93.33